OpenStudy (anonymous):

csc -pi/2 exact value

OpenStudy (shamil98):

\[\csc = \frac{ 1 }{ \sin \Theta }\]

OpenStudy (anonymous):



OpenStudy (anonymous):

Is that right?

OpenStudy (shamil98):

The exact result after plugging into a calculator is 1.

OpenStudy (anonymous):

I'll tell you how you figure that out though :P

OpenStudy (shamil98):

You could go on the unit circle

OpenStudy (shamil98):

Find where sin (cos,sin) is -pi/2

OpenStudy (shamil98):

and put into 1/sin form.

OpenStudy (anonymous):

Lol was just drawing it, but yeah >_>

OpenStudy (anonymous):

1/1 = 1

OpenStudy (shamil98):

:D

OpenStudy (anonymous):

I mean 1/ -1 = -1 not 1 >_>

OpenStudy (anonymous):

Since it's -pi/2

OpenStudy (shamil98):

is it? o.o , oh my bad, the typos of today -_-